OpenStudy (anonymous):

y= ln(e^(x^2-1) find y'(1) a. 0 b. 1/2 c. 1 d. 2 e. undefined please help and show all steps thank you!

myininaya (myininaya):

\[\text{ if } y=\ln(e^{g(x)}) \text{ then} y=g(x)\ln(e)=g(x)(1)=g(x) =>y'=g'(x)\]